Given the functions f(x)= x^2 - 9x - 3 and g(x)= x - 3, find f(g(x))
Answer by edjones(8007)

f(x)= x^2 - 9x - 3 and g(x)= x - 3
(x-3)^2-9(x-3)-3
x^2-6x+9-9x+27-3
f(g(x))=x^2-15x+33